OpenStudy (anonymous):

The graph of a polynomial passes through points A(2,5), B(3,2), and C(5,-1) use Gaussian elimination to find the polynomial. what is the best way to start?

OpenStudy (anonymous):

HELP!!

OpenStudy (amistre64):

3 points, 3 unknowns ...

OpenStudy (amistre64):

2^2 2 1 5 3^2 3 1 2 5^2 5 1 -1 4 2 1 5 9 3 1 2 25 5 1 -1 row reduce to your hearts content

OpenStudy (anonymous):

Ok thank you for the matrices but how did you get to that first step?

OpenStudy (amistre64):

hmm, well we need to be able to fit these point to some curve. 2 points define a line, 3 points define a quadratic ... we need some quadratic equation that satisfies:\[ax^2+bx+c=y\]for the stated points \[2^2a+2b+1c=5\]\[3^2a+3b+1c=2\]\[5^2a+5b+1c=-1\] then its just a matter of pulling off the coefficents to form the matrix

OpenStudy (anonymous):

That makes sense thank you very much! :)

OpenStudy (amistre64):

youre welcome :)